# Stabilizer Ranks, Barnes Wall Lattices and Magic Monotones

### v1

@@ 6 Exponential Exact Stabilizer Rank for a Product State
The idea of using these $V_{k}$ spaces originates from their use in Mehraban and Tahmasbi’s proof of the quadratic lower bound[7]. However, it turns out these are the very objects that even Lovitz and Steffan [8] used in their proof of these results, but instead treated them as algebraic structures called varieties. Their proof technique also directly gives us the result that the set of maximal rank states has full measure. However, , as we directly make use of the notion of distances, we believe our techniques might be more suitable for working on the problem of approximate stabilizer ranks (see Problem 56).
@@ 7 Directions for Future work
Problem 56.
Find examples of $\delta(n)$ for which $\chi_{\delta_{n}}(\mathinner{|{\psi}\rangle})=2^{\Omega(n)}$ for some product state $\mathinner{|{\psi}\rangle}$.
@@
Finally, it can be useful generalize the notions of the monotones $\mathcal{N}$ and $\mathcal{N}_{\delta}$ using the Barnes Wall lattices corresponding to $\mathbb{Z}[e^{i\pi/4}]$, as defined in [12]. This is because then there would be some vector proportional to $\mathinner{|{T}\rangle}^{\otimes n}$ on the lattice, which is not the case for $BW_{n}$. Using that, we can get a corresponding lower bound on the $T$-counts of state preparation analogous to Theorem 5.
@@
To counteract this, instead of directly working with the $l_{2}$ norm of the vectors, we can work with some norm function $F$ applied to $\mathinner{\langle{\psi|\psi}\rangle}$. The norm function defined in [12] is a candidate, but since it isn’t multiplicative, most of the arguments in Section 4 fail.
Next, we can consider the field norm for the extension $\mathbb{Q}(\sqrt{2})/\mathbb{Q}$. Since it can be negative, we may further take the absolute value. Using this norm, almost all the arguments in Section 4 can be recreated, except, we have no knowledge about the minimal vectors with respect to this norm. It would still be the case that the stabilizer states have norm 1, but the arguments in [12] do not exclude the possibility of even smaller vectors with respect to this norm. In case there are indeed even smaller vectors, then this would lose any significance as a measure of magic.
Problem 57.
@@ 8 Acknowledgments
ARK and PS are very grateful to David Gosset for suggesting the connection between Barnes Wall Lattices and Stabilizer Ranks and for subsequent related discussions. We also thank Graeme Smith, Benjamin Lovitz and Yash Totani for discussions.
PS is supported by Mike and Ophelia Lazaridis Fellowship. ARK thanks NTT research and NSERC for financial support. Research at Perimeter Institute and IQC is supported by the Government of Canada through Innovation, Science and Economic Development Canada, and by the Province of Ontario through the Ministry of Research, Innovation and Science.

### v2

@@ 6 Exponential Exact Stabilizer Rank for a Product State
The idea of using these $V_{k}$ spaces originates from their use in Mehraban and Tahmasbi’s proof of the quadratic lower bound[Saeed]. However, it turns out these are the very objects that even Lovitz and Steffan [Lovitz2022newtechniques] used in their proof of these results, but instead treated them as algebraic structures called varieties. Their proof technique also directly gives us the result that the set of maximal rank states has full measure. However, as we directly make use of the notion of distances, we believe our techniques might be more suitable for working on the problem of approximate stabilizer ranks (see Problem 59).
@@ 7 Directions for Future work
Problem 59.
Find examples of $\delta(n)$ for which $\chi_{\delta(n)}(\mathinner{|{\psi}\rangle})=2^{\Omega(n)}$ for some product state $\mathinner{|{\psi}\rangle}$.
@@
Finally, it can be useful generalize the notions of the monotones $\mathcal{N}$ and $\mathcal{N}_{\delta}$ using the Barnes Wall lattices corresponding to $\mathbb{Z}[e^{i\pi/4}]$, as defined in [kliuchnikov2024stabilizer]. This is because then there would be some vector proportional to $\mathinner{|{T}\rangle}^{\otimes n}$ on the lattice, which is not the case for $BW_{n}$. Using that, we can get a corresponding lower bound on the $T$-counts of state preparation analogous to Theorem 6.
@@
To counteract this, instead of directly working with the $l_{2}$ norm of the vectors, we can work with some norm function $F$ applied to $\mathinner{\langle{\psi|\psi}\rangle}$. The norm function defined in [kliuchnikov2024stabilizer] is a candidate, but since it isn’t multiplicative, most of the arguments in Section 4 fail.
Next, we can consider the field norm for the extension $\mathbb{Q}(\sqrt{2})/\mathbb{Q}$. Since it can be negative, we may further take the absolute value. Using this norm, almost all the arguments in Section 4 can be recreated, except, we have no knowledge about the minimal vectors with respect to this norm. It would still be the case that the stabilizer states have norm 1, but the arguments in [kliuchnikov2024stabilizer] do not exclude the possibility of even smaller vectors with respect to this norm. In case there are indeed even smaller vectors, then this would lose any significance as a measure of magic.
Problem 60.
@@ 8 Acknowledgments
ARK and PS are very grateful to David Gosset for suggesting the connection between Barnes Wall Lattices and Stabilizer Ranks and for subsequent related discussions. We also thank Graeme Smith, Benjamin Lovitz and Yash Totani for discussions. PS is supported by Mike and Ophelia Lazaridis Fellowship. ARK was supported in part by Canada’s NSERC, NTT Research, and the Royal Bank of Canada. IQC and the Perimeter Institute (PI) are supported in part by the Government of Canada through ISED and the Province of Ontario.